OpenStudy (lin.ivory):

simplify: 4-8(sin^2 6x) please show steps!! thank you!

OpenStudy (ash2326):

\[4-8 (\sin^2 6x)\] Hi @lin.ivory Do you know the formula of sin 3x?

OpenStudy (lin.ivory):

i know "sin2x" but ive never heard of sin3x

OpenStudy (ash2326):

Ok do you have options? Cause it can be simplified in several ways

OpenStudy (lin.ivory):

a. cos12x b. 2cos6 c. 4cos6x d. 4cos12x the answer is D, but i dont know how to get there! =(

OpenStudy (ash2326):

Ok got it:D You know what's cos 2x?

OpenStudy (lin.ivory):

OHHHHHHHHHHHHHHHHHHH!!!!!!!!!!!!! so do i factor out the 4 and get 4(cos2(6x)) and then just multiply to get 4(cos12x)??!!! AWESOMENESS!! THANK YOU!!

OpenStudy (ash2326):

\[\mathbb{You're\ welcome:)}\] \[ \huge \mathbb{And\ Welcome\ To\ Open\ Study}\]